What are the solutions for the equation (5x-2)^2=10
Answer by richwmiller(17219)

(5x-2)^2=10
get square root of both sides
|5x-2|=+\-sqrt(10)
5x=2+\-sqrt(10)
5x=2+sqrt(10)
5x=2-sqrt(10)
divide both sides by 5
x=(2-sqrt(10))/5=1/5((2-sqrt(10))
x=(2+sqrt(10))/5=1/5((2+sqrt(10))